OpenStudy (gabyramirezzz):

Algebra 2 Word Problem

OpenStudy (gabyramirezzz):

A bank has loaned out $500,000 part of it at 6% annum and the rest at 11% annum. If the bank receives $43,750 in interest each year how much was loaned at 6%?

OpenStudy (agent0smith):

If x is how much in the 6% and y at 11%: x + y = 500,000 since the total is $500K 0.06x + 0.11y = 43,750 since the total interest is $43,750

OpenStudy (gabyramirezzz):

oh okay

OpenStudy (gabyramirezzz):

so where do we go from here?

OpenStudy (agent0smith):

Use elimination or substitution method.

OpenStudy (gabyramirezzz):

so would the next step be .06(500000)=43750?

OpenStudy (agent0smith):

No, the next step is to solve the system of equations, using elimination or substitution x + y = 500,000 0.06x + 0.11y = 43,750

OpenStudy (gabyramirezzz):

oh okay so it would be 0.06x + 0.11(500000) = 43750?

OpenStudy (gabyramirezzz):

@DanJS can you help?

OpenStudy (danjs):

yeah , what part you not getting?

OpenStudy (gabyramirezzz):

when i calculate it, it comes out to -11250, but I thought it couldn't be a negative

OpenStudy (danjs):

Take the two equations... x + y = 500,000 0.06x + 0.11y = 43,750 Solve the first for either x or y. y = 500000 - x Put that value for y into the second equation 0.06x + 0.11*(500000 - x) = 43750 solve for x

OpenStudy (danjs):

Looks like you put in 5000000 for y instead of 500000 - x

OpenStudy (gabyramirezzz):

yeah, i definitely messed up. thank you.

OpenStudy (danjs):

Or you can multiply the first one by 11 and the second by 100 -- 11x + 11y = 500000 6x + 11y = 43750 subtract the second equation from the first equation 5x = 456250

OpenStudy (danjs):

welcome